OpenStudy (anonymous):

Solve the equation. (x-3)(x+6)=10

OpenStudy (mathstudent55):

First, multiply out the left side using FOIL. (x - 3)(x + 6) = 10 x^2 + 6x - 3x - 18 = 10 Now subtract 10 from both sides and collect like therms. x^2 + 6x - 3x - 18 - 10 = 0 x^2 + 3x - 28 = 0 Now try factoring the left side. Are you able to continue now?

OpenStudy (anonymous):

is it -7,4?

OpenStudy (mathstudent55):

That's the final solution. This is how you get there: (x +7)(x -4) = 0 x + 7 = 0 or x - 4 = 0 x = -7 or x = 4

OpenStudy (anonymous):

thank you !

OpenStudy (mathstudent55):

wlcm